# Multiplier modules of Hilbert C*-modules revisited

### v2

@@ Abstract.
The theory of multiplier modules of Hilbert C*-modules is reconsidered to obtain more properties of these special Hilbert C*-modules. The property of a Hilbert C*-module to be a multiplier C*-module is shown to be an invariant with respect to the consideration as a left or right Hilbert C*-module in the sense of a imprimitivity bimodule in strong Morita equivalence theory. The interrelation of the C*-algebras of ”compact” operators, the Banach algebras of bounded module operators and the Banach spaces of bounded module operators of a Hilbert C*-module to its C*-dual Banach C*-module, are characterized for pairs of Hilbert C*-modules and their respective multiplier modules. The structures on the latter are always isometrically embedded into the respective structures on the former. Examples for which continuation of these kinds of bounded module operators from the initial Hilbert C*-module to its multiplier module fails are given, however existing continuations turn out to be always unique. Similarly, bounded modular functionals from both kinds of Hilbert C*-modules to their respective C*-algebras of coefficients are compared, and eventually existing continuations are shown to be unique.
@@
Considering multiplier modules as full left Hilbert C*-modules $X$ over a C*-algebra $A$ we obtain that they are at the same time full right Hilbert C*-modules and multiplier modules over the respective C*-algebras ${\rm K}_{A}(X)$. This reminds special imprimitivity bimodules, however no new type of Morita equivalence can be derived. Considering pairs $(X,M(X))$ of Hilbert C*-modules and their multiplier modules we consider the interrelations of comparable types of operator algebras of bounded module operators over them and continuation problems of operators from $X$ to $M(X)$, the same for bounded modular functionals. So, for ”compact” operator algebras ${\rm K}_{A}(X)$ is *-isometrically embedded into $K_{M(A)}(M(X))$, but the former might be smaller than the latter one. Furthermore, the Banach algebra of bounded module operators ${\rm End}_{M(A)}(M(X))$ is isometrically embedded into ${\rm End}_{A}(X)$, as well as the Banach space ${\rm End}_{M(A)}(M(X),M(X)^{\prime})$ into ${\rm End}_{A}(X,X^{\prime})$, and again the former might be smaller than the latter one. Consequently, not any operator of the smaller structures on $X$ can be continued to a bounded operator on the larger structures $M(X)$ obeying strict convergence. If such a continuation exists it is unique. The same picture can be obtained for bounded modular functionals on $M(X)$ and on $X$, respectively, so we found examples of pairs of Hilbert C*-modules $(X,M(X))$ with $X^{\bot}=\{0\}$ for which no general Hahn-Banach type theorem can be obtained. Remarkably, there does not exist any non-trivial bounded module map from $M(X)$ to $M(A)$ vanishing on $X$.
@@ 1. Introduction
We denote C*-algebras by $A,B$. In case a C*-algebra $A$ is non-unital, general C*-multiplier theory provides us with some derived structures like multiplier algebras $M(A)$, left and right multiplier algebras $LM(A)$ and $RM(A)$, resp., and quasi-multiplier spaces $QM(A)$. To calculate these linear spaces any faithful *-representation of $A$ on a Hilbert space $H$ can be used. The calculation environment is the von Neumann algebra generated by the faithfully *-represented C*-algebra $A$, or $B(H)$ itself, cf. [35, Ch. 3], [39]. The unital C*-algebra $M(A)$ is defined as
@@
Inventing different kinds of strict topologies intrinsic characterizations of these structures as certain topological completions of $A$ on bounded sets of $A$ are available. For comprehensive sources we refer to the book by P. Ara and M. Matthieu [2] and to [32, 33, 7, 16]. Note, that multiplier algebras might admit an entire lattice of non-unital, two-sided, non-isomorphic ideals $A_{\alpha}$ such that $M(A_{\alpha})=M(A_{\beta})$ for any two of them, cf. [25]. Also, either $M(A)=LM(A)$ and $M(A)=QM(A)$ at the same time, or $LM(A)$ is strictly larger than $M(A)$ and $QM(A)$ is strictly larger than $LM(A)$, cf. [8, Cor. 4.18].
@@
We would like to consider Hilbert C*-modules over (non-unital, in general) C*-algebras and, in particular, their multiplier modules and related structures. By convention all Hilbert C*-modules are right C*-modules, at the first glance. However, for full Hilbert $A$-modules one can obtain an operator-valued inner product turning them into full left Hilbert ${\rm K}_{A}(X)$-modules, and vice versa. So, the point of view decides which of the two Hilbert C*-module structures is primary and which is secondary. This kind of consideration will be used frequently in the present paper. The interested reader is refered to the theory of strong Morita equivalence of C*-algebras and to C*-correspondences or imprimitivity bimodules, see [41, 14]. We use the term ’imprimitivity bimodule’ instead of the newer term ’$A$-$B$ Hilbert bimodule’ and avoid the elder term ’C*-correspondence’ due to changes of its exact meaning in the recent literature.
@@
A pre-Hilbert $A$-module over a C*-algebra $A$ is an $A$-module $X$ equipped with an $A$-valued, non-degenerate mapping .,.$\langle.,.\rangle:X\times X\to A$ being conjugate-$A$-linear in the first argument and $A$-linear in the second one, and satisfying $\langle x,x\rangle\geq 0$ for every $x\in X$. The map .,$\langle$.,$\rangle$ . , . ⟩ is called the $A$-valued inner product on $X$. A pre-Hilbert $A$-module $\{X,\langle.,.\rangle\}$ is Hilbert if and only if it is complete with respect to the norm .=.,.$\|.\|=\|\langle.,.\rangle\|^{1/2}_{A}$. We always assume that the complex linear structures of $A$ and $X$ are compatible. A Hilbert $A$-module $\{X,\langle.,.\rangle\}$ over a C*-algebra $A$ is full if the norm-closed $A$-linear hull $\langle X,X\rangle$ of the range of the inner product coincides with $A$. Two (full) Hilbert $A$-modules $\{X,\langle.,.\rangle_{X}\}$ and $\{Y,\langle.,.\rangle_{Y}\}$ over a fixed C*-algebra $A$ are unitarily equivalent (or unitarily isomorphic) iff there exists a bounded invertible adjointable map $T:X\to Y$ such that .,.$\langle.,.\rangle_{X}=\langle T(.),T(.)\rangle_{Y}$ on $X$. The $A$-dual Banach $A$-module $X^{\prime}$ of a Hilbert $A$-module $X$ is defined as the set of all bounded $A$-linear maps from $X$ into $A$. It might not be a Hilbert $A$-module itself, cf. [16]. But, $X$ is always canonically isometrically embedded into $X^{\prime}$ as a Banach $A$-submodule via the identification of $x\in X$ with $\langle x,\cdot\rangle\in X^{\prime}$. Note, that two $A$-valued inner products on a Hilbert $A$-module $X$ inducing equivalent norms on $X$ might not be unitarily isomorphic, cf. [8, 16]. Thus, full Hilbert C*-modules are always a triple of the C*-algebra $A$ of coefficients, the Banach $A$-module $X$ and the $A$-valued inner product on $X$. We omit the explicit reference to the C*-valued inner product in places where its definition formula is standard or only its existence is important.
@@ 2. On multiplier modules
Let us show a way to define multiplier modules $M(X)$ of given (full) Hilbert $A$-modules $X$ as a certain related Hilbert $M(A)$-module, and let us provide some crucial properties of them. We start with a class of extensions of a given full Hilbert $A$-module $X$ as defined in [5]:
@@
    ${\rm Im}(\Phi)=YA=\{za:z\in Y,a\in A\}=\{x\in Y:\langle x,x\rangle\in A\}$ (by the Hewitt-Cohen factorization theorem, [40, Thm. 4.1], [41, Prop. 2.31], [24, Thm. 23.22]).
@@
Note, that $\Phi$ is an $A$-linear isometry of Hilbert $A$-modules and, hence, in case of a surjective mapping a unitary map, preserving $A$-valued inner products up to unitary equivalence, cf. [30], [15, Thm. 5], [44, Thm. 1.1]. So, $Y$ and $\Phi(Y)$ are unitarily equivalent Hilbert C*-modules. In the sequel we consider the C*-algebras $A$ and $B$ as a Hilbert $A$-module and as a Hilbert $B$-module over itself, respectively, setting $\langle a,b\rangle=a^{*}b$ for any two C*-algebra elements $a,b$. Then $A$ and $B$ are *-isometrically isomorphic to the C*-algebras ${\rm K}_{A}(A)$ and ${\rm K}_{B}(B)$, respectively. By [47, Prop. 2.2.16] these *-isomorphisms extend to *-isomorphisms of $M(A)$ and of $M(B)$ with the C*-algebras ${\rm End}_{M(A)}^{*}(A)$ and ${\rm End}_{M(B)}^{*}(B)$, respectively. We shall use these identifications freely.
@@
In [5] $M(X)$ is shown to be an essential extension of $X$ identifing $X$ and ${\rm K}_{A}(A,X)$ as the subset $\{za:z\in M(X),a\in A\}$ isometrically. In fact, $M(X)$ is the largest essential extension of $X$, because $M(A)$ is the largest essential extension of $A$ containing *-isomorphic copies of all C*-algebras $B$ which contain *-isomorphic copies of $A$ as an essential ideal, [5, Thm. 1.2]. This justifies the point of view on $M(X)$ as a Hilbert C*-module version of the multiplier algebra in C*-theory. For a given Hilbert $A$-module $\{X,\langle.,.\rangle_{X}\}$ the respective multiplier module $M(X)$ is unique up to unitary isomorphism of Hilbert $M(A)$-modules.
Let us explain the definition of multiplier modules $M(X)$ for non-full Hilbert $A$-modules $X$. $\langle X,X\rangle$ is a two-sided norm-closed ideal of $A$. If $X$ is a non-full Hilbert $A$-module form the derived full Hilbert $A$-module $X_{c}$ by adding a copy of $A$ as an orthogonal direct summand to $X$, $X_{c}=X\oplus A$, .,.$\langle.,.\rangle_{c}=\langle.,.\rangle_{X}+\langle.,.\rangle_{A}$. Then construct its multiplier module $M(X_{c})$. Using the mapping picture of Definition 2.2 of $M(X_{c})$ we see that $M(X_{1}\oplus X_{2})=M(X_{1})\oplus M(X_{2})$ using modular projection operators onto each of the orthogonal summands. So, we have an orthogonal decomposition $M(X_{c})=M(X)\oplus M(A)$, taking the first orthogonal direct summand and Hilbert $M(A)$-module as the definition of the multiplier module $M(X)$ of $X$. Clearly, the (non-unital, in general) C*-algebra $\langle X,X\rangle$ is a two-sided norm-closed ideal of $A$, and the C*-algebra $\langle M(X),M(X)\rangle$ is a two-sided norm-closed ideal of $M(A)$ containing $\langle X,X\rangle$. However, the latter might be non-unital, so we should be more careful with their C*-algebraic interrelations. In general, $\langle X,X\rangle$ is a two-sided norm-closed ideal in $\langle M(X),M(X)\rangle$. The uniqueness results for the pairings $(X,M(X))$ are preserved. It is important to realize, that $M(X)$ might depend on the choice of the C*-algebra acting on $X$, e.g. for $X=A$ with $A$ a non-unital C*-algebra we have $M(A_{A})={\rm End}_{A}^{*}(A,A)=M(A)$, but $M(A_{M(A)})={\rm End}_{M(A)}^{*}(M(A),A)=A$. Last but not least, if the C*-algebra $A$ is non-unital then for $X=l_{2}(A)$ the multiplier module $M(l_{2}(A))$ equals to
@@
– a self-dual Hilbert W*-module over $M(A)=B(l_{2})$, because $A$ is stable, cf. [5, Ex. 2.2] and [37].
@@
The first assertion follows from (iv) of Definition 2.1 combining $\Phi_{2}$ with $\Phi_{1}^{-1}$ with respect to the set $\{za:z\in M(X),a\in A\}$ in the largest essential extension $M(X)$, cf. [42, Prop. 1.4]. The second assertion can be illustrated by example, however $M(X_{1})A_{1}$ is obviously not isometrically isomorphic to $M(X_{2})A_{2}$ by supposition.
@@
By [42, Prop. 1.1] both $X_{1}$ and $X_{2}$ admit a canonical isometric modular embedding into $M(X_{1})$ and into $M(X_{2})$, respectively, as Hilbert $M(A_{1})$-submodules. However, the multiplier module of $A_{2}$ with respect to $A_{1}$, $M_{A_{1}}(A_{2})$, is a non-unital C*-algebra.
@@
For $X,Y$ Hilbert $A$-modules each operator $T\in{\rm End}_{A}^{*}(X,Y)$ has an extension $T_{M}\in{\rm End}_{M(A)}^{*}(M(X),M(Y))$ of the same norm value obtained as the strict continuation of $T$. Therefore, it is uniquely determined. Moreover, every operator in ${\rm End}_{M(A)}^{*}(M(X),M(Y))$ arises this way, i.e. the C*-algebras ${\rm End}_{M(A)}^{*}(M(X),M(Y))$ and ${\rm End}_{A}^{*}(X,Y)$ are *-isomorphic, [5, Thm. 2.3]. Since a full Hilbert $A$-module $X$ is not only a right Hilbert $A$-module, but also a full left Hilbert ${\rm K}_{A}(X)$-module by the theory of strong Morita equivalence of the C*-algebras $A$ and ${\rm K}_{A}(X)$, $X$ can be considered as a C*-correspondence or as an imprimitivity bimodule of these two C*-algebras. By [28] the C*-algebra ${\rm End}_{A}^{*}(X)$ can be considered as the multiplier algebra of the C*-algebra ${\rm K}_{A}(X)$. Consequently, $M(X)$ is a full left Hilbert $M({\rm K}_{A}(X))$-module. So, for (full) Hilbert ${\rm K}_{A}(X)$-modules $X$ the Hilbert $M({\rm K}_{A}(X))$-module $M(X)$ can be identified with the (left) multiplier module of $X$ w.r.t. the pairing $({\rm K}_{A}(X),M({\rm K}_{A}(X)))$, too. This makes the property of a (full) Hilbert C*-module to be a multiplier module invariant under the choice of the point of view as a (full) left or right Hilbert C*-module. For similar thoughts compare with [5, pp. 20-21, (a)-(b)]

### v3

@@ Abstract.
The theory of multiplier modules of Hilbert C*-modules is reconsidered to obtain more properties of these special Hilbert C*-modules. The property of a Hilbert C*-module to be a multiplier C*-module is shown to be an invariant with respect to the consideration as a left or right Hilbert C*-module in the sense of a imprimitivity bimodule in strong Morita equivalence theory. The interrelation of the C*-algebras of ”compact” operators, the Banach algebras of bounded module operators and the Banach spaces of bounded module operators of a Hilbert C*-module to its C*-dual Banach C*-module, are characterized for pairs of Hilbert C*-modules and their respective multiplier modules. The structures on the latter are always isometrically embedded into the respective structures on the former. Examples are given for which continuation of these kinds of bounded module operators from the initial Hilbert C*-module to its multiplier module fails. However, existing continuations turn out to be always unique. Similarly, bounded modular functionals from both kinds of Hilbert C*-modules to their respective C*-algebras of coefficients are compared, and eventually existing continuations are shown to be unique.
@@
Considering multiplier modules as full left Hilbert C*-modules $X$ over a C*-algebra $A$ we obtain that they are at the same time full right Hilbert C*-modules and multiplier modules over the respective C*-algebras ${\rm K}_{A}(X)$. This reminds special imprimitivity bimodules, however no new type of Morita equivalence can be derived. Considering pairs $(X,M(X))$ of Hilbert C*-modules and their multiplier modules we consider the interrelations of comparable types of operator algebras of bounded module operators over them and continuation problems of operators from $X$ to $M(X)$, the same for bounded modular functionals. So, for ”compact” operator algebras ${\rm K}_{A}(X)$ is *-isometrically embedded into $K_{M(A)}(M(X))$, but the former might be smaller than the latter one. Furthermore, the Banach algebra of bounded module operators ${\rm End}_{M(A)}(M(X))$ is isometrically embedded into ${\rm End}_{A}(X)$, as well as the Banach space ${\rm End}_{M(A)}(M(X),M(X)^{\prime})$ into ${\rm End}_{A}(X,X^{\prime})$, and again the former might be smaller than the latter one. Consequently, not any operator of the smaller spaces on $X$ can be continued to a bounded operator on the larger spaces on $M(X)$ obeying strict convergence. If such a continuation exists it is unique. The same picture can be obtained for bounded modular functionals on $M(X)$ and on $X$, respectively, so we found examples of pairs of Hilbert C*-modules $(X,M(X))$ with $X^{\bot}=\{0\}$ for which no general Hahn-Banach type theorem can be obtained. Remarkably, there does not exist any non-trivial bounded module map from $M(X)$ to $M(A)$ vanishing on $X$.
@@ 1. Introduction
We denote C*-algebras by $A,B$. In case a C*-algebra $A$ is non-unital, general C*-multiplier theory provides us with some derived structures like multiplier algebras $M(A)$, left and right multiplier algebras $LM(A)$ and $RM(A)$, resp., and quasi-multiplier spaces $QM(A)$. To calculate these linear spaces any faithful *-representation of $A$ on a Hilbert space $H$ can be used. The calculation environment is the von Neumann algebra generated by the faithfully *-represented C*-algebra $A$, or $B(H)$ itself, cf. [38, Ch. 3], [42]. The unital C*-algebra $M(A)$ is defined as
@@
Inventing different kinds of strict topologies intrinsic characterizations of these structures as certain topological completions of $A$ on bounded sets of $A$ are available. For comprehensive sources we refer to the book by P. Ara and M. Matthieu [2] and to [35, 36, 8, 18]. Note, that multiplier algebras might admit an entire lattice of non-unital, two-sided, non-isomorphic ideals $A_{\alpha}$ such that $M(A_{\alpha})=M(A_{\beta})$ for any two of them, cf. [27]. Also, either $M(A)=LM(A)$ and $M(A)=QM(A)$ at the same time, or $LM(A)$ is strictly larger than $M(A)$ and $QM(A)$ is strictly larger than $LM(A)$, cf. [9, Cor. 4.18].
@@
We would like to consider Hilbert C*-modules over (non-unital, in general) C*-algebras and, in particular, their multiplier modules and related structures. By convention all Hilbert C*-modules are right C*-modules, at the first glance. However, for full Hilbert $A$-modules one can obtain an operator-valued inner product turning them into full left Hilbert ${\rm K}_{A}(X)$-modules, and vice versa. So, the point of view decides which of the two Hilbert C*-module structures is primary and which is secondary. This kind of consideration will be used frequently in the present paper.
@@
A pre-Hilbert $A$-module over a C*-algebra $A$ is an $A$-module $X$ equipped with an $A$-valued, non-degenerate mapping $\langle.,.\rangle:X\times X\to A$ being conjugate-$A$-linear in the first argument and $A$-linear in the second one, and satisfying $\langle x,x\rangle\geq 0$ for every $x\in X$. The map $\langle.,.\rangle$ is called the $A$-valued inner product on $X$. A pre-Hilbert $A$-module $\{X,\langle.,.\rangle\}$ is Hilbert if and only if it is complete with respect to the norm $\|.\|=\|\langle.,.\rangle\|^{1/2}_{A}$. We always assume that the complex linear structures of $A$ and $X$ are compatible. A Hilbert $A$-module $\{X,\langle.,.\rangle\}$ over a C*-algebra $A$ is full if the norm-closed $A$-linear hull $\langle X,X\rangle$ of the range of the inner product coincides with $A$. Two (full) Hilbert $A$-modules $\{X,\langle.,.\rangle_{X}\}$ and $\{Y,\langle.,.\rangle_{Y}\}$ over a fixed C*-algebra $A$ are unitarily equivalent (or unitarily isomorphic) iff there exists a bounded invertible adjointable map $T:X\to Y$ such that $\langle.,.\rangle_{X}=\langle T(.),T(.)\rangle_{Y}$ on $X$. The $A$-dual Banach $A$-module $X^{\prime}$ of a Hilbert $A$-module $X$ is defined as the set of all bounded $A$-linear maps from $X$ into $A$. It might not be a Hilbert $A$-module itself, cf. [18]. But, $X$ is always canonically isometrically embedded into $X^{\prime}$ as a Banach $A$-submodule via the identification of $x\in X$ with $\langle x,\cdot\rangle\in X^{\prime}$. Note, that two $A$-valued inner products on a Hilbert $A$-module $X$ inducing equivalent norms on $X$ might not be unitarily isomorphic, cf. [9, 18]. Thus, full Hilbert C*-modules are always a triple of the C*-algebra $A$ of coefficients, the Banach $A$-module $X$ and the $A$-valued inner product on $X$. We omit the explicit reference to the C*-valued inner product in situations where its definition formula is standard or only its existence is important.
@@
We are partially interested in strong Morita equivalence of C*-algebras and in imprimitivity bimodules realizing those equivalences. A C*-correspondence from a C*-algebra $B$ to a C*-algebra $A$ (or an $B$-AA C*-correspondence) is a Hilbert module $X$ over $A$ together with a nondegenerate *-homomorphism from $B$ to ${\rm End}^{*}_{A}(X)$, which introduces a left $B$-module structure on $X$. For any $B$-AA C*-correspondence $X$ the actions extend uniquely to multiplier algebras of $B$ and of $A$ so that we may always treat $X$ as a $M(B)$-$M(A)$ C*-correspondence. To relate it to the multiplier $M(B)$-$M(A)$ bimodule $M(X)$, let us note that whenever $Y$ is a Hilbert $A$-module, then ${\rm End}^{*}_{A}(Y,X)$ is naturally a C*-correspondence from $M(B)$ to ${\rm End}^{*}_{A}(Y)$. In particular, ${\rm End}^{*}_{A}(Y,X)$ contains ${\rm K}_{A}(Y,X)$ as a sub-C*-correspondence. An $B$-AA imprimitivity bimodule is a full $B$-AA C*-correspondence $X$ which is also a full left Hilbert $B$-module and the two $B$- resp. $A$-valued inner products satisfy the equality $\langle x,y\rangle_{B}z=x\langle y,z\rangle_{A}$ for an $x,y,z\in X$. In particular, for imprimitivity bimodules $X$, $X$ is a full $B$-AA C*-correspondence such that the left action $\phi:B\to{\rm End}^{*}_{A}(X)$ restricts to a *-isomorphism of $B$ with ${\rm K}_{A}(X)$. Two C*-algebras $B$ and $A$ are said to be strongly Morita equivalent iff there exists an imprimitivity bimodule (or a $B$-AA imprimitivity bimodule) connecting them. It is a real equivalence relation, so one can form classes of strongly Morita equivalent C*-algebras sharing important properties among their elements. The interested reader is referred to [44, 16, 13, 14, 12, 15] to learn more about the theory of strong Morita equivalence of C*-algebras and about C*-correspondences or imprimitivity bimodules. If $X$ and $Y$ are, respectively, $B$-AA and $C$-$B$ imprimitivity bimodules, then their algebraic tensor product $X\odot_{B}Y$ is a $C$-AA imprimitivity bimodule, and one can define two inner products on $X\odot_{B}Y$ by
$\langle x\otimes z,y\otimes t\rangle_{A}=\langle\langle y,x\rangle_{B}z,t\rangle_{A}\,,\,\,\langle x\otimes z,y\otimes t\rangle_{C}=\langle x,y\langle t,z\rangle_{B}\rangle_{C}$
for any $x,y\in X$, $z,t\in Y$. This way $X\odot_{B}Y$ becomes an $C$-AA imprimitivity bimodule. Remarkably, the set of all $B$-AA imprimitivity bimodules connecting two given strongly Morita equivalent C*-algebras $A$ and $B$ can be quite manifold. Using the defined tensor product operation the problem burns down to the consideration of the Picard group ${\rm Pic}(A)$ of one C*-algebra $A$ of unitary isomorphism classes of AA-AA imprimitivity bimodules for an arbitrarily selected element $A$ of a given class of strongly Morita eqivalent to $A$ C*-algebras. Note, that the inverse of an imprimitivity bimodule $X$ in ${\rm Pic}(A)$ is its dual module $\tilde{X}$}. If $Y$ is an $B$-AA imprimitivity bimodule, then the map $X\to{\tilde{Y}}\otimes_{A}X\otimes_{A}Y$ induces an isomorphism of ${\rm Pic}(A)$ and ${\rm Pic}(B)$, and [7, Theorem 1.2] shows that the Picard groups are stable isomorphism invariant. For a non-trivial example cf. [31] for the Picard groups of irrational rotation algebras. For a comprehensive account see [11, sec. 6-7].
@@ 2. On multiplier modules
Let us show a way to define multiplier modules $M(X)$ of given (full) Hilbert $A$-modules $X$ as a certain related Hilbert $M(A)$-module, and let us provide some important properties of them. We start with a class of extensions of a given full Hilbert $A$-module $X$ as defined in [5]:
@@
    ${\rm Im}(\Phi)=YA=\{za:z\in Y,a\in A\}=\{x\in Y:\langle x,x\rangle\in A\}$ (by the Hewitt-Cohen factorization theorem, [43, Thm. 4.1], [44, Prop. 2.31], [26, Thm. 23.22]).
@@
Note, that $\Phi$ is an $A$-linear isometry of Hilbert $A$-modules and, hence, in case of a surjective mapping a unitary map, preserving $A$-valued inner products up to unitary equivalence, cf. [33], [17, Thm. 5], [47, Thm. 1.1]. So, $Y$ and $\Phi(Y)$ are unitarily equivalent Hilbert C*-modules. In the sequel we consider the C*-algebras $A$ and $B$ as a Hilbert $A$-module and as a Hilbert $B$-module over itself, respectively, setting $\langle a,b\rangle=a^{*}b$ for any two C*-algebra elements $a,b$. Then $A$ and $B$ are *-isometrically isomorphic to the C*-algebras ${\rm K}_{A}(A)$ and ${\rm K}_{B}(B)$, respectively. By [50, Prop. 2.2.16] these *-isomorphisms extend to *-isomorphisms of $M(A)$ and of $M(B)$ with the C*-algebras ${\rm End}_{M(A)}^{*}(A)$ and ${\rm End}_{M(B)}^{*}(B)$, respectively. We shall use these identifications freely.
@@
In [5] $M(X)$ is shown to be an essential extension of $X$ identifying $X$ and ${\rm K}_{A}(A,X)$ as the subset $\{za:z\in M(X),a\in A\}$ isometrically. In fact, $M(X)$ is the largest essential extension of $X$, because $M(A)$ is the largest essential extension of $A$ containing *-isomorphic copies of all C*-algebras $B$ which contain *-isomorphic copies of $A$ as an essential ideal, [5, Thm. 1.2]. This justifies the point of view on $M(X)$ as a Hilbert C*-module analog of the multiplier algebra in C*-theory. For a given Hilbert $A$-module $\{X,\langle.,.\rangle_{X}\}$ the respective multiplier module $M(X)$ is unique up to unitary isomorphism of Hilbert $M(A)$-modules.
Let us explain the definition of multiplier modules $M(X)$ for non-full Hilbert $A$-modules $X$. $\langle X,X\rangle$ is a two-sided norm-closed ideal of $A$. If $X$ is a non-full Hilbert $A$-module form the derived full Hilbert $A$-module $X_{c}$ by adding a copy of $A$ as an orthogonal direct summand to $X$, $X_{c}=X\oplus A$, $\langle.,.\rangle_{c}=\langle.,.\rangle_{X}+\langle.,.\rangle_{A}$. Then construct its multiplier module $M(X_{c})$. Using the mapping picture of Definition 2.2 of $M(X_{c})$ we see that $M(X_{1}\oplus X_{2})=M(X_{1})\oplus M(X_{2})$ using modular projection operators onto each of the orthogonal summands. So, we have an orthogonal decomposition $M(X_{c})=M(X)\oplus M(A)$, taking the first orthogonal direct summand and Hilbert $M(A)$-module as the definition of the multiplier module $M(X)$ of $X$. Clearly, the (non-unital, in general) C*-algebra $\langle X,X\rangle$ is a two-sided norm-closed ideal of $A$, and the C*-algebra $\langle M(X),M(X)\rangle$ is a two-sided norm-closed ideal of $M(A)$ containing $\langle X,X\rangle$. However, the latter might be non-unital, so we should be more careful with their C*-algebraic interrelations. In general, $\langle X,X\rangle$ is a two-sided norm-closed ideal in $\langle M(X),M(X)\rangle$. The uniqueness results for the pairings $(X,M(X))$ are preserved. It is important to realize, that $M(X)$ might depend on the choice of the C*-algebra acting on $X$, e.g. for $X=A$ with $A$ a non-unital C*-algebra we have $M_{A}(A)={\rm End}_{A}^{*}(A,A)=M(A)$, but $M_{M(A)}(A)={\rm End}_{M(A)}^{*}(M(A),A)=A$. Last but not least, if the C*-algebra $A$ is non-unital then for $X=l_{2}(A)$ the multiplier module $M(l_{2}(A))$ equals to
@@
– a self-dual Hilbert W*-module over $M(A)=B(l_{2})$, because $A=K(l_{2})$ is stable, cf. [5, Ex. 2.2] and [40].
@@
The first assertion follows from (iv) of Definition 2.1 combining $\Phi_{2}$ with $\Phi_{1}^{-1}$ with respect to the set $\{za:z\in M(X),a\in A\}$ in the largest essential extension $M(X)$, cf. [45, Prop. 1.4]. The second assertion can be illustrated by example (see below), however $M(X_{1})A_{1}$ is obviously not isometrically isomorphic to $M(X_{2})A_{2}$ by supposition.
@@
By [45, Prop. 1.1] both $X_{1}$ and $X_{2}$ admit a canonical isometric modular embedding into $M(X_{1})$ and into $M(X_{2})$, respectively, as Hilbert $M(A_{1})$-submodules. However, the multiplier module of $A_{2}$ with respect to $A_{1}$, $M_{A_{1}}(A_{2})$, is a non-unital C*-algebra.
@@
For $X,Y$ Hilbert $A$-modules each operator $T\in{\rm End}_{A}^{*}(X,Y)$ has an extension $T_{M}\in{\rm End}_{M(A)}^{*}(M(X),M(Y))$ of the same norm value obtained as the strict continuation of $T$. Therefore, it is uniquely determined. Moreover, every operator in ${\rm End}_{M(A)}^{*}(M(X),M(Y))$ arises this way, i.e. the C*-algebras ${\rm End}_{M(A)}^{*}(M(X),M(Y))$ and ${\rm End}_{A}^{*}(X,Y)$ are *-isomorphic, [5, Thm. 2.3]. Since a full Hilbert $A$-module $X$ is not only a right Hilbert $A$-module, but also a full left Hilbert ${\rm K}_{A}(X)$-module by the theory of strong Morita equivalence of the C*-algebras $A$ and ${\rm K}_{A}(X)$, $X$ can be considered as a C*-correspondence or as an imprimitivity bimodule of these two C*-algebras. By [30] the C*-algebra ${\rm End}_{A}^{*}(X)$ can be considered as the multiplier algebra of the C*-algebra ${\rm K}_{A}(X)$. Consequently, $M(X)$ is a full left Hilbert $M({\rm K}_{A}(X))$-module. So, for (full) Hilbert ${\rm K}_{A}(X)$-modules $X$ the Hilbert $M({\rm K}_{A}(X))$-module $M(X)$ can be identified with the (left) multiplier module of $X$ w.r.t. the pairing $({\rm K}_{A}(X),M({\rm K}_{A}(X)))$, too. This makes the property of a (full) Hilbert C*-module to be a multiplier module invariant under the choice of the point of view as a (full) left or right Hilbert C*-module. For similar thoughts compare with [5, pp. 20-21, (a)-(b)], [16, Props. 1.3, 1.6, Remark].